OpenStudy (anonymous):

Triangle ABC has vertices A (1,5) B (9,7)and C (7,3). M is the midpoint of AB and N is the midpoint of BC. Which of the following is the length of MN? A. 3 B. 4 C. 5 D. 6

OpenStudy (anonymous):

First: Use the midpoint formula to find M and N. \(\huge (\frac{ x_2+x_1}{ 2 },\frac{ y_2+y_1 }{ 2 })\)

OpenStudy (anonymous):

M is the midpoint of A and B, and A (1, 5) & B (9, 7). x1 y1 x2 y2

OpenStudy (anonymous):

N is the midpoint of B and C, and A (9, 7) & B (7, 3). x1 y1 x2 y2

OpenStudy (anonymous):

Once you found the points of M and N, use the distance formula to find the length of MN. DISTANCE FORMULA: \[\huge \sqrt{(x_2 - x_1)^2+(y_2-y_1)^2}\]